OpenStudy (jamaljunior):

The formula for the circumference of a circle is C = 2πr. Use this formula to write a formula for the RADIUS of a circle.

OpenStudy (jamaljunior):

did not mean to close T_T

jhonyy9 (jhonyy9):

yes this is the formula for circumference of a circle C = 2pir so how you get from this isolate the r ?

jhonyy9 (jhonyy9):

ok. than you divide both sides by 2pi what will get ?

OpenStudy (jamaljunior):

idk

jhonyy9 (jhonyy9):

why ? C/2pi = 2pir / 2pi C/2pi = ?

OpenStudy (jamaljunior):

what is pir

jhonyy9 (jhonyy9):

pi is 3,14 and r is the radius

jhonyy9 (jhonyy9):

pi*r

OpenStudy (jamaljunior):

what is the radius

jhonyy9 (jhonyy9):

in your question of above wrote exercise there is RADIUS so this r is the radius noted

jhonyy9 (jhonyy9):

r is the radius of this circle with circumference C = 2pi*r

jhonyy9 (jhonyy9):

ok. ?

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (jamaljunior):

yes

jhonyy9 (jhonyy9):

ok . was my pleasure good luck bye bye

jhonyy9 (jhonyy9):

or do you have questions ?

OpenStudy (jamaljunior):

what answer?

jhonyy9 (jhonyy9):

what is the answer here ? so given C = 2pi*r than you divide both sides by 2pi will get C/2pi = 2pi*r/2pi C/2pi = ?

jhonyy9 (jhonyy9):

do you see there on the right side will simplify the 2pi from numerator and denominator and will get the r what you need getting for right answer ok. ?

OpenStudy (jamaljunior):

ok

jhonyy9 (jhonyy9):

so than now is clearly sure for you - yes ?

OpenStudy (jamaljunior):

yes ty

jhonyy9 (jhonyy9):

np yw anytime bye bye